Part 1 - A certificate of deposit often charges a penalty for withdrawing funds before the maturity date. If the penalty involves two months of interest, what would be the amount for early withdrawal on a CD paying 4 percent and valued at $26,000? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Part 2 - A payday loan company charges 5 percent interest for a four-week period. What would be the annual interest rate from that company? (Assume an even 52 weeks per year. Enter your answer as a percent rounded to 1 decimal place.)
Answer #1

1.
=Value of CD*rate*penalty months/12
=26000*4%*2/12
=173.33

2.
APR=5%*52/4=65.00000%

EAR=(1+5%)^(13)-1=88.56491%